3x+y=10
y=-2x+4
solve this by using substitution

Answer by robertb(5830)

Substitute the bottom equation into the top equation:
3x+(-2x+4)=10.
3x-2x+4 = 10.
x = 6. Then
y = -2(6) + 4 = -12+4 = -8.